# Itô’s formula for flows of conditional measures on semimartingales

### v2

@@ Abstract.
Motivated by recent development of mean-field systems with common noise, this paper establishes Itô’s formula for flows of conditional probability measures under a common filtration associated with general semimartingales. This generalizes existing works on flows of conditional measures on Itô processes and flows of deterministic measure on general semimartingales. The key technical components involve constructing conditional independent copies and establishing the equivalence between stochastic integrals with respect to the conditional law of semimartingales and the conditional expectation of stochastic integrals with respect to copies of semimartingales. Itô’s formula is then established for cylindrical functions through conditional independent copies, and extended to the general case through function approximations.
@@ 1. Introduction
The classical Itô’s formula for semimartingales is generally viewed as the stochastic counterpart of the chain rule in calculus. It stands as one of the key cornerstones of stochastic analysis [29], establishing an intrinsic connection between partial differential equations (PDEs) and diffusion processes. This formula plays a critical role in the theory of stochastic control.
With the recent development of theory of mean-field games and mean-field controls [6, 8, 11], Itô’s formula has been extended to the flows of probability measures. Initially applied to diffusions [6], this extension was further developed for (discontinuous) semimartingales [24]. The generalized Itô’s formula has been instrumental for deriving the master equations in mean-field games and the Bellman dynamic programming equation for McKean–Vlasov control problems, as discussed in [11] and the references therein.
One of the most recent advances in mean-field theory have been concerning mean-field controls and games with common noise (see for instance [1, 2, 3, 5, 9, 10, 13, 15, 18, 23, 27, 30, 31, 32, 33, 35, 40]) or with stochastic processes involving jumps ([7, 22, 25, 26, 28, 36]). In either case, Itô’s formula has been developed respectively in [11] and in [24]. In order to analyze rigorously the McKean–Vlasov dynamics with both semimartingales and idiosyncratic and common noises, it is necessary to establish the Itô’s formula for flows of conditional laws with semimartingales, which does not seem to exist to the best of our knowledge. This is the primary focus of this paper.
@@
This paper establishes Itô’s formula (Theorem 2.7) for flow of conditional laws $\mu_{t}=$Law$(X_{t}|{\mathcal{G}}_{t})$ driven by (possible discontinuous) semimartingales $\{X_{t}\}_{t\in[0,T]}$ with a general form of common noise represented by a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$ and ${\mathcal{G}}={\mathcal{G}}_{T}$.
The primary task to analyze the Itô integral $\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}^{\mathcal{G}}$ with $\{X_{t}^{\mathcal{G}}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]\}_{t\in[0,T]}$ over some appropriate adapted process $\{\eta_{t}\}_{t\in[0,T]}$. This is crucial as the integral of the conditional expectation such as $\int_{0}^{t}\eta_{s}{\mathrm{d}}X^{\mathcal{G}}_{s}$ is not necessarily equivalent to the conditional expectation of the integral of the form ${\mathbb{E}}\big{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}|{\mathcal{G}}\big{]}$ (See discussions and examples in Section 3). Our analysis is built on several key technical components: constructing conditional independent copies of general stochastic processes (Theorem 3.1), establishing the equivalence between the conditional expectation of stochastic integrals with respect to seminmartingales and stochastic integrals with respect to the conditional law of semimartingales, and building the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditional independent copies (Theorem 3.3).
With these key technical components, Itô’s formula for flows of conditional laws on semimartingales is established (Theorem 2.7): it is first derived for cylindrical functions (i.e., smooth mean-field functions with integrable forms 4.1)), which are then shown to be dense in the desired function space; finally, localization argument along with proper form of dominating convergence theorem finishes the task.
@@
Conditional independent copied has been treated informally and in certain forms in the existing literature, for instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise in [11]. In this paper, we have generalized the construction of conditional independent copies for a broad class of semimartingales and with an arbitrary sub $\sigma$-algebra. This concept may be of independent theoretical interest, besides its crucial role in our derivation of Itô’s formula for flow of conditional laws on semimartingales.
There are several methods for deriving Itô’s formula and its variants, including the time discretization approach in [6] and [36] for mean-field jump diffusions, the density approach in [8] using the Fokker–Planck equation, and the particle approximation approach [11, 14], [19, 38] to approximate flows of measures by flows of empirical measures. The cylindrical function approach has proved appropriate and powerful for analysis of general semimartingales. They are initially explored in Fleming–Viot processes [21], later adopted in the analysis of polynomial diffusions [17], and most recently used in Itô’s formula for flow of measures with (discontinuous) semimartingales [24].
Meanwhile, linear derivatives on the space of probability measures are known to be appropriate for characterizing the infinitesimal changes in the functional of controlled McKean–Vlasov processes when jumps are added; this has been noted previously in [7] and further explored in [24] and [39].
@@
  • ${\mathbb{N}}^{+}$ denotes the the set of all possible natural numbers, and we fix $d\in{\mathbb{N}}^{+}$.
@@
  • $C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with continuous derivatives up to the $k$-th order, and $C_{b}^{k}({\mathbb{R}}^{d})\subset C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with bounded and continuous derivatives up to the $k$-th order. And we use the convention $C({\mathbb{R}}^{d})=C^{0}({\mathbb{R}}^{d})$ and $C_{b}({\mathbb{R}}^{d})=C^{0}_{b}({\mathbb{R}}^{d})$.
  • For a random variable $X$, $\|X\|_{L^{p}}$ represents its $p$-th moment, i.e., $\|X\|_{L^{p}}={\mathbb{E}}[|X|^{p}]^{\frac{1}{p}}$.
  • For vectors $\bm{a}=(a_{i})_{i=1,..,d},\bm{b}=(b_{i})_{i=1,..,d}\in{\mathbb{R}}^{d}$, we denote $\bm{a}\cdot\bm{b}=\sum_{i=1}^{d}a_{i}b_{i}$. For matrices $\bm{C}=(C_{ij})_{i,j=1,..,d},\bm{D}=(D_{ij})_{i,j=1,..,d}\in{\mathbb{R}}^{d\times d}$, we denote $\bm{C}:\bm{D}=\sum_{i,j=1}^{d}C_{ij}D_{ij}$.
  • For $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ and $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ such that $\int_{{\mathbb{R}}^{d}}|\varphi(x)|{\mathrm{d}}\mu(x)<\infty$, we set
    $\langle\mu,\varphi\rangle:=\int_{{\mathbb{R}}^{d}}\varphi(x){\mathrm{d}}\mu(x).$
@@ 2. Assumptions and main result
Given two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$- semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ for $T>0$ and $d\in\mathbb{N}^{+}$ on a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and $\mu=\big{(}\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})\big{)}_{t\in[0,T]}$ with a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$. Our focus is to develop an Itô’s formula for the functional $\Phi(\mu_{t},\bm{Y}_{t})$ with the function $\Phi:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, under suitable regularity conditions.
2.1. Assumptions
Throughout the paper, we will make the following assumptions on the semimartingales $\bm{X}$ and $\bm{Y}$, the filtrations ${\mathbb{G}}$ and ${\mathbb{F}}$, and the function $\Phi$. These are standard assumptions to ensure that stochastic integrals are properly defined (see [37]) with the presence of common noise (see [34]) and for flow of measures on semimartingales (see [24, 16]).
Assumptions on semimartingales.
Assumption 2.1 (Semimartingale).
  • We assume $\bm{X},\bm{Y}\in\mathcal{H}^{p}$ for $1\leq p\leq\infty$ [37, Section V.2], with
    $\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}^{p}}:=\inf_{\bm{M},\bm{V}}\bigg{\|}|\bm{X}_{0}|+\sqrt{[\bm{M},\bm{M}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{V}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$ (2.1)
    $\big{\|}\bm{Y}\big{\|}_{{\mathcal{H}}^{p}}:=\inf_{\bm{N},\bm{U}}\bigg{\|}|\bm{Y}_{0}|+\sqrt{[\bm{N},\bm{N}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{U}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$
@@ 2.1. Assumptions
    $\bm{X}_{t}=\bm{X}_{0}+\bm{M}_{t}+\bm{V}_{t}\text{ and }\bm{Y}_{t}=\bm{Y}_{0}+\bm{N}_{t}+\bm{U}_{t}.$
    Here $\bm{V}=(\bm{V}_{t})_{t\in[0,T]}$ and $\bm{U}=(\bm{U}_{t})_{t\in[0,T]}$ are adapted càdlàg processes of finite variation with $\bm{V}_{0}=\bm{U}_{0}=\bm{0}$, and $\bm{M}=(\bm{M}_{t})_{t\in[0,T]}$ and $\bm{N}=(\bm{N}_{t})_{t\in[0,T]}$ are ${\mathbb{F}}$-local martingales such that $\bm{M}_{0}=\bm{N}_{0}=\bm{0}$.
  • We assume also
@@
Under these assumptions, a proper form of dominated convergence theorem has been established ([37, Page 273, Lemma]) and the following two propositions hold.
Proposition 2.2 ([37, Section V, Theorem 2]).
For any $1\leq p\leq\infty$, consider a norm on the space of ${\mathbb{R}}$-valued adapted càdlàg processes (or the space of ${\mathbb{R}}$-valued adapted càglàd processes) defined by
@@
Then, for $1\leq p<\infty$, there exists a constant $c_{p}$ depending only on $p$ such that
@@
Proposition 2.3 (Emery inequality, [37, Section V, Theorem 3]).
Let $Z$ be an ${\mathbb{R}}$-valued semimartingale, $H$ is an ${\mathbb{R}}$-valued adapted càglàd process, and $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ ($1\leq p,q\leq\infty)$, then
$\Big{\|}\int_{0}^{\cdot}H_{s}{\mathrm{d}}Z_{s}\Big{\|}_{{\mathcal{H}}^{r}}\leq\|H\|_{{\mathcal{S}}^{p}}\|Z\|_{{\mathcal{H}}^{q}}.$
@@
These propositions are important for the subsequent analysis, for instance in the localization argument for establishing Itô’s lemma for the flow of conditional measure on semimartinagles, as will be clear in Section 4.2.
Assumptions on filtration.
Assumption 2.5 (Filtration).
@@
Sub-filtration ${\mathbb{G}}\subset{\mathbb{F}}$ satisfies the compatibility assumption or the conditional independence condition. That is, for each $0\leq t\leq T$, ${\mathcal{F}}_{t}$ and ${\mathcal{G}}_{T}$ are independent, given ${\mathcal{G}}_{t}$, written as ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t}.$
@@
In the context of mean field games, this assumption ensures a weak closure of adapted processes, as first noted in [12, Lemma 3.11], and then stated more explicitly in [4, Theorem 5.4].
One can deduce from is assumption that $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, for any ${\mathbb{F}}$-adapted process $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ almost surely for each $t\leq T$, according to a special form of the (H) hypothesis by freezing ${\mathcal{G}}_{u}={\mathcal{G}}_{T}$ when $u\geq T$ ([20]). This allows us to define a cádlág version of $\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})$ by taking $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, which will be assumed throughout the paper.
Assumptions on function $\Phi$. We will adopt the the linear derivative adapted from [16], which has shown to be appropriate for the analyzing the flow of measures on semimartingales [24, 16].
@@
Definition 2.6 ($C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ functions).

### v3

@@ Abstract.
Motivated by recent developments of mean-field systems with common noise, this paper establishes Itô’s formula for flows of conditional probability measures under a common filtration associated with general semimartingales. It generalizes existing works on flows of conditional measures on Itô processes and flows of deterministic measures on general semimartingales. It is derived via the cylindrical function approach. One key technical component involves constructing conditionally independent copies of stochastic processes, which allows for establishing the equivalence between stochastic integrals with respect to the conditional laws of semimartingales and the conditional expectation of stochastic integrals with respect to copies of semimartingales.
As an application, a class of control problems for jump-diffusions with partial observations is analyzed.
@@ 1. Introduction
The classical Itô’s formula for semimartingales, the stochastic equivalent of the chain rule in calculus, is a cornerstone of stochastic analysis [36]. It intrinsically links partial differential equations (PDEs) with diffusion processes. Recent advancements in mean-field game and control theories [9, 11, 14] have led to its extension to flows of probability measures, initially for diffusions [9] and later for discontinuous semimartingales ([31, 53]). This generalized Itô’s formula is crucial for deriving master equations in mean-field games and Bellman dynamic programming equations for McKean–Vlasov control problems, see [14] and the references therein.
Most recent studies in mean-field theory involve controls and games with common noise (see for instance [1, 2, 5, 7, 12, 13, 17, 19, 24, 30, 34, 40, 41, 44, 45, 48, 54]) or stochastic jump processes ([10, 28, 32, 33, 35, 49]). In either case, Itô’s formula has been developed respectively in ([9, 14]) and ([31, 53]). In order to analyze rigorously the McKean–Vlasov dynamics with both semimartingales and idiosyncratic and common noises, it is necessary to establish the Itô’s formula for flows of conditional laws with semimartingales, which does not seem to exist, to the best of our knowledge. This is the primary focus of this paper.
@@
[…]
One key step is to analyze the Itô integral $\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}^{\mathcal{G}}$ with $\{X_{t}^{\mathcal{G}}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]\}_{t\in[0,T]}$ over some appropriate adapted process $\{\eta_{t}\}_{t\in[0,T]}$. This is crucial as the integral of the conditional expectation such as $\int_{0}^{t}\eta_{s}{\mathrm{d}}X^{\mathcal{G}}_{s}$ is not necessarily equivalent to the conditional expectation of the integral of the form ${\mathbb{E}}\big{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}|{\mathcal{G}}\big{]}$ (See discussions and examples in Section 3). Our analysis is built on two key technical components: one is to construct conditionally independent copies of general stochastic processes (Theorem 3.1), the other is to establish the equivalence between the conditional expectation of stochastic integrals with respect to semimartingales and stochastic integrals with respect to the conditional law of semimartingales, as well as the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditionally independent copies (Theorem 3.5).
With these key technical components, Itô’s formula for flows of conditional laws on semimartingales is established (Theorem 2.11): the result is first established for cylindrical functions, defined as smooth mean-field functions with integrable representations (see Definition 4.1). Since these functions are dense in the desired function space, the general form is then obtained by a localization argument combined with suitable versions of the dominated convergence theorem.
This Itô’s formula (Theorem 2.11) is applied to analyze control problems with partially observed jump-diffusions. An appropriate form of the verification theorem 6.2 is derived, which recovers the results in [4] as special cases.
@@
One of the key technical elements is the construction of conditionally independent copies of stochastic processes. This concept has been treated informally, for instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise in [14].
Our construction of conditionally independent copies of general stochastic processes is inspired by the proofs of Yamada-Watanabe theorem (see for instance [39, 42]) and the theory of weak and strong solutions of stochastic differential equations (see for instance [37, 43]). Most existing methods specify the common information $\mathcal{G}$} as a stochastic input, and make use of the existence of the regular conditional distribution to construct an extended probability space. In contrast, we construct conditionally independent copies directly, rather than utilizing the regular conditional distribution. (See the discussion at the beginning of Section 5 for a detailed comparison).
Our construction approach generalizes the construction of conditionally independent copies to a broad class of semimartingales and allows for an arbitrary sub $\sigma$-algebra. This result may be of independent theoretical interest, besides its essential role in our derivation of Itô’s formula for flow of conditional laws on semimartingales.
A crucial assumption in our analysis is a condition on the filtration known as the compatibility condition. This condition appears in various contexts under different names, such as immersion [38], very good extensions [37], and natural extensions [50]. In the theory of filtration enlargement, it is known as the H-hypothesis ([8, 26]), which plays a central role for characterizing semimartingale properties with filtration expansion or shrinkage. Following [43], we adopt the term compatibility to be aligned with recent literature on mean-field games ([6, 14, 15, 16, 46]).
Finally, it is worth pointing out that there are several methods to derive Itô’s formula and its variants, including the time discretization approach in [9] and [49] for mean-field jump diffusions, the density approach in [11] using the Fokker–Planck equation, and the particle approximation approach of [14, 18], [25, 52] by flows of empirical measures. The cylindrical function approach has proved appropriate and powerful for analysis of general semimartingales. It is initially explored for Fleming–Viot processes [27], later adopted in the analysis of polynomial diffusions [23], and for Itô’s formulas on flow of measures with (discontinuous) semimartingales [31] and for non-anticipative maps of càdlàg rough paths [22].
@@
  • We use the bold letter $\bm{x},\bm{X}$ to denote a (random) vector with multiple dimensions or a matrix depending on the context. $|\bm{X}|$ denotes the Euclidean norm of a vector $\bm{X}\in{\mathbb{R}}^{d}$, or the Frobenius norm of a matrix.
@@
  • $C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with continuous derivatives up to the $k$-th order, and $C_{b}^{k}({\mathbb{R}}^{d})\subset C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with bounded and continuous derivatives up to the $k$-th order, with the convention $C({\mathbb{R}}^{d})=C^{0}({\mathbb{R}}^{d})$ and $C_{b}({\mathbb{R}}^{d})=C^{0}_{b}({\mathbb{R}}^{d})$.
  • For a $d$-dimensional vector-valued random variable $\bm{X}$, $\|\bm{X}\|_{L^{p}}$ represents its $p$-th moment, i.e., $\|\bm{X}\|_{L^{p}}={\mathbb{E}}[|\bm{X}|^{p}]^{\frac{1}{p}}$.
  • For vectors $\bm{a}=(a_{i})_{i=1,..,d},\bm{b}=(b_{i})_{i=1,..,d}\in{\mathbb{R}}^{d}$, denote $\bm{a}\cdot\bm{b}=\sum_{i=1}^{d}a_{i}b_{i}$. For matrices $\bm{C}=(C_{ij})_{i,j=1,..,d},\bm{D}=(D_{ij})_{i,j=1,..,d}\in{\mathbb{R}}^{d\times d}$, denote $\bm{C}:\bm{D}=\sum_{i,j=1}^{d}C_{ij}D_{ij}$.
  • For $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ and $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ such that $\int_{{\mathbb{R}}^{d}}|\varphi(\bm{x})|{\mathrm{d}}\mu(\bm{x})<\infty$, set
    $\langle\mu,\varphi\rangle:=\int_{{\mathbb{R}}^{d}}\varphi(\bm{x}){\mathrm{d}}\mu(\bm{x}).$
  • For a càdlàg process $\bm{X}_{t}$, we adopt the standard notation:
    $\bm{X}_{t-}:=\lim_{s\nearrow t}\bm{X}_{s},\quad\Delta\bm{X}_{t}:=\bm{X}_{t}-\bm{X}_{t-},$
    and $\bm{X}^{c}$ to represent the continuous part of the process.
@@ 2. Assumptions and main result
Fix $T>0$ and $d\in\mathbb{N}^{+}$. Consider two ${\mathbb{R}}^{d}$-valued $\mathbb{F}$}-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ on a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and $\mu=\big{\{}\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})\big{\}}_{t\in[0,T]}$ for a sub-filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$. Our focus is to establish an Itô’s formula for the functional $\Phi(\mu_{t},\bm{Y}_{t})$, with $\Phi:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ for some $p\in[2,\infty)$ under suitable regularity conditions.
2.1. Assumptions and preliminary tools
Throughout the paper, we will make the following assumptions on the semimartingales $\bm{X}$ and $\bm{Y}$, the filtrations $\mathbb{G}$} and $\mathbb{F}$}, and the function $\Phi$. These are standard assumptions to ensure that stochastic integrals are properly defined ([51]) with the presence of common noise ([47]), and for flow of measures on semimartingales ([21, 31]).
First, we specify the assumptions on the semimartingales $\bm{X}$ and $\bm{Y}$.
Assumption 2.1.
For some $1\leq p\leq\infty$,
  • $\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}^{p}}=\inf_{\bm{M},\bm{V}}\bigg{\|}|\bm{X}_{0}|+\sqrt{[\bm{M},\bm{M}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{V}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$ (2.1)
    $\big{\|}\bm{Y}\big{\|}_{{\mathcal{H}}^{p}}=\inf_{\bm{N},\bm{U}}\bigg{\|}|\bm{Y}_{0}|+\sqrt{[\bm{N},\bm{N}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{U}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$
@@ 2.1. Assumptions and preliminary tools
    $\bm{X}_{t}=\bm{X}_{0}+\bm{M}_{t}+\bm{V}_{t}\text{ and }\bm{Y}_{t}=\bm{Y}_{0}+\bm{N}_{t}+\bm{U}_{t},$
    where $\bm{V}=(\bm{V}_{t})_{t\in[0,T]}$ and $\bm{U}=(\bm{U}_{t})_{t\in[0,T]}$ are adapted càdlàg processes of finite variation with $\bm{V}_{0}=\bm{U}_{0}=\bm{0}$, and $\bm{M}=(\bm{M}_{t})_{t\in[0,T]}$ and $\bm{N}=(\bm{N}_{t})_{t\in[0,T]}$ are $\mathbb{F}$}-local martingales such that $\bm{M}_{0}=\bm{N}_{0}=\bm{0}$.
@@
There are two propositions that are necessary for the subsequent analysis, especially for the localization argument in establishing the Itô’s lemma for the flow of conditional measures on semimartingales, as will be clear in Section 4.2.
Proposition 2.2 ([51, Section V, Theorem 2]).
For any $1\leq p\leq\infty$, define a norm on the space $D$ of all $\mathbb{R}$}-valued adapted càdlàg processes, such that for $H=(H_{t})_{t\in[0,T]}\in D$,
@@
Then, there exists a constant $c_{p}$ depending only on $p$ such that
@@
Proposition 2.3 (Emery inequality, [51, Section V, Theorem 3]).
Let $Z$ be an $\mathbb{R}$}-valued semimartingale, $H$ be an $\mathbb{R}$}-valued adapted càdlàg process, and $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ ($1\leq p,q\leq\infty)$, then
$\Big{\|}\int_{0}^{\cdot}H_{s-}{\mathrm{d}}Z_{s}\Big{\|}_{{\mathcal{H}}^{r}}\leq\|H\|_{{\mathcal{S}}^{p}}\|Z\|_{{\mathcal{H}}^{q}}.$
@@
Moreover, a proper form of dominated convergence theorem has been established as follows.
Lemma 2.5.
([51, Page 273, Lemma]) Let $p,q,r$ be given such that $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ where $1<r<+\infty$. Let $Z$ be a semimartingale such that $\|Z\|_{\mathcal{H}_{q}}<+\infty$, and let $\{H^{n}=(H^{n}_{t})_{t\in[0,T]}\}_{n\in{\mathbb{N}}}$ and $Y=(Y_{t})_{t\in[0,T]}$ be semimartingales such that $\|H^{n}\|_{{\mathcal{S}}^{p}}<+\infty$ for all $n\in{\mathbb{N}}$, $\|Y\|_{{\mathcal{S}}^{p}}<+\infty$ and $|H^{n}_{t}|\leq Y_{t}$. Suppose $\lim_{n\to\infty}H^{n}_{t-}(\omega)=0$, for all $(t,\omega)$. Then
$\lim_{n\to\infty}\bigg{\|}\int_{0}^{T}H^{n}_{s-}{\mathrm{d}}Z_{s}\bigg{\|}_{{\mathcal{H}}^{r}}=0.$
Next, we impose the compatibility condition for filtrations, in the following sense.
Assumption 2.6 (Compatibility).
@@
The sub-filtration ${\mathbb{G}}\subset{\mathbb{F}}$ satisfies the compatibility condition, if for each $0\leq t\leq T$, ${\mathcal{F}}_{t}$ and ${\mathcal{G}}_{T}$ are conditionally independent given ${\mathcal{G}}_{t}$, expressed as ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t}.$
Remark 2.7.
There are a number of alternative characterizations for the compatibility condition, such as the immersion or the H-hypothesis, in the literature of filtration enlargement and shrinkage (see for instance [8, 26, 38]). Given two filtrations ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$ and ${\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]}$ defined on the same space, with ${\mathbb{G}}\subset{\mathbb{F}}$, the following are known to be equivalent ([8, Theorem 3]):
  • ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, for all $0\leq t\leq T.$ (Compatibility condition).
  • Every square integrable $\mathbb{G}$}-martingale is an $\mathbb{F}$}-martingale. (H-hypothesis).
  • ${\mathbb{E}}[Z|{\mathcal{G}}_{t}]$ = ${\mathbb{E}}[Z|{\mathcal{F}}_{t}]$ a.s., for each $t\in[0,T]$ and each bounded ${\mathcal{F}}_{T}$-measurable random variable $Z$.
@@
In the context of mean-field games, this compatibility assumption ensures a weak closure of adapted processes, as first noted in [15, Lemma 3.11] and stated more explicitly in [6, Theorem 5.4].
Moreover, as shown in the following lemma, this compatibility condition allows us to define a càdlàg version of $\mu_{t}$ by taking $\mu_{t}=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, which will be used throughout the paper.
Lemma 2.8.
Under Assumption 2.6, $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})$ almost surely.
Proof.
It suffices to show that for any bounded ${\mathcal{G}}_{T}$- measurable random variable $A$, ${\mathbb{E}}\big{[}\varphi(\bm{X}_{t})A\big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(\bm{X}_{t})|{\mathcal{G}}_{t}\big{]}A\Big{]}.$ This is guaranteed by ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$ and
${\mathbb{E}}\big{[}\varphi(X_{t})A\big{]}$ $={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})A|{\mathcal{G}}_{t}\big{]}\Big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}{\mathbb{E}}\big{[}A|{\mathcal{G}}_{t}\big{]}\Big{]}$
$={\mathbb{E}}\Big{[}{\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}A\big{|}{\mathcal{G}}_{t}\Big{]}\Big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}A\Big{]}.$
@@
Finally, we will assume $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ according to the following definition.
Definition 2.9 ($C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$).